OpenStudy (anonymous):

Solve the system by substitution: -4.5x - 2y = - 12.5 3.25x - y =0.75

OpenStudy (anonymous):

easier if you multipiy all by 100 and start with \[-450x-200y=-1250\\ 325x-100y=758\]

OpenStudy (anonymous):

damn typo \[-450x-200y=-1250\\ 325x-100y=75\]

OpenStudy (anonymous):

multiply the second equation by \(-2\) and get \[-450-200y=-1250\\ -650x+200y=-150\]

OpenStudy (anonymous):

add and you should be in good shape

OpenStudy (anonymous):

What am I adding? Sorry, I'm just not sure how to complete the problem :P